# Recovery of rational functions via Hankel pencil method and sensitivities of the poles

### v1

@@ 1 Introduction
We study reconstruction of rational functions (1.1) using as input data a finite set of their Fourier coefficients. We use the property that Fourier coefficients of rational functions have the exponential structure and we reconstruct this exponential structure in the frequency domain applying the ESPRIT method [17, 15]. The idea of recovery of rational functions by estimation of signal parameters in the frequency domain is not new. Similar technique for the recovery of or approximation by rational functions was employed in [6, 18, 19, 20]. We present modification of these ideas, by applying the ESPRIT method for the exponential recovery. Since ESPRIT is based on using the Hankel pencil, it makes the method suitable for sensitivity analysis.
The other methods for the recovery of rational functions (in time domain) such as the AAA method [13], the RKFIT (Rational Krylov Fitting) method [3], the Vector Fitting algorithm [12] and the Loewner framework method [1], were considered in [11]. Note that, in [6, 18, 19, 20], similar to [11], reconstruction from given noisy data is limited to numerical observations.
@@
In [21], for the Loewner framework method the authors study sensitivities of Loewner pencil eigenvalues (which are poles of a transfer function defined by the corresponding Loewner matrices) with respect to unstructured and structured perturbations. Following approach in [21], we present the sensitivity analysis for the poles computed by our method. According to our method, poles located inside and outside the unit circle are reconstructed separately as eigenvalues of some Hankel matrix pencils. Further we consider two kinds of Hankel pencil perturbations. For the sensitivity $\rho$ with respect to unstructured perturbations, we present the connection with condition numbers of Vandermonde matrices that are used in factorizations of the corresponding Hankel matrices. The sensitivity $\eta$ with respect to structured perturbations of Hankel pencil is studied in [22]. We present modification of these results for our special reconstruction problem. Note that the sensitivity of eigenvalues of Hankel pencil perturbations was studied also in [10], [2]. In [10] and [2], the authors measure both structured and unstructured perturbations by the same quantity. Following approach in [21], we consider two different quantities for structured and unstructured sensitivities respectively. Further using sensitivity analysis, we explain reconstruction errors for poles from noisy Fourier coefficients. We refer also to the paper [4] for further application of the sensitivity analysis of Hankel pencil.
@@
The method offered in this paper is also very much related to our ESPIRA method for the exponential recovery [7][9]. In ESPIRA, we use the property that the classical Fourier coefficients of exponential sums have a rational structure and to reconstruct this rational structure we apply the AAA method for rational approximation [13] in the frequency domain. Thus the idea of the method for rational recovery used in this paper can be considered as converse to the idea of our ESPIRA method.
@@
Outline. In Section 2, we present the main ideas of the ESPRIT method for exponential recovery. In Section 3 we consider sensitivity analysis for Hankel pencil eigenvalues. Subsection 3.1 is dedicated to unstructured perturbations and in Subsection 3.2, we consider structured perturbations. In Section 4, we describe our main results. In Subsection 4.1, we present main ideas of our method for the recovery of rational functions (1.1). Subsection 4.2 is dedicated to the sensitivity analysis of poles reconstructed with our method. Finally, in Section 5, we present numerical experiments.
@@
Notation. Let us agree that throughout the paper, we will use the matrix notation $\mathbf{X}_{L,K}:=(X_{ij})_{i,j=1}^{L,K}$ for complex matrices of size $L\times K$ and the submatrix notation $\mathbf{X}_{L,K}(m:n,k:\ell)$, to denote a submatrix of $\mathbf{X}_{L,K}$ with rows indexed $m$ to $n$ and columns indexed $k$ to $\ell$, where first row and first column has index 1 (even though the row and column indices for the definition of the matrix may start with 0). For square matrices, we often use the short notation $\mathbf{X}_{K}$ instead of $\mathbf{X}_{K,K}$. By $\left(\mathbf{X}_{L,K}\right)^{\dagger}$ we denote the Moore-Penrose inverse of $\mathbf{X}_{L,K}$. We consider $\|\mathbf{X}_{L,K}\|_{2}:=\sigma_{\max}(\mathbf{X}_{L,K})$ the 2-norm or spectral norm, and $\|\mathbf{X}_{L,K}\|_{\mathrm{F}}:=\left(\sum\limits_{i=1}^{L}\sum\limits_{j=1}^{K}|X_{ij}|^{2}\right)^{1/2}=\left(\sum\limits_{i=1}^{\min\{L,K\}}\sigma_{i}^{2}\right)^{1/2}$ the Frobenius norm of the matrix $\mathbf{X}_{L,K}$. The $\ell_{1}$- and $\ell_{2}$-norm of the vector $\boldsymbol{x}=(x_{1},\ldots,x_{K})$ are given as usually by $\|\boldsymbol{x}\|_{1}=\sum\limits_{j=1}^{K}|x_{j}|$ and $\|\boldsymbol{x}\|_{2}=\left(\sum\limits_{j=1}^{K}|x_{j}|^{2}\right)^{1/2}$ respectively.
@@ 3 Sensitivities of Hankel pencil eigenvalues
In this section, we consider sensitivities of eigenvalues of the Hankel pencil (2.3). The concept of sensitivity arises from the perturbation of eigenvalues of the matrix pencil and for Hankel matrices was introduced in [10] and further considered in [2]. In [21], the notion of sensitivities was defined for Loewner matrix pencil. In the paper [21], the authors define two sensitivities, which correspond to structured and unstructured perturbations. Using similar ideas, in [22], the author considered the structured sensitivity of Hankel matrix pencil.
@@ 3.1 Sensitivity of Hankel pencil eigenvalues with unstructured perturbations
$(\boldsymbol{p}^{T}+(\boldsymbol{p}^{(1)})^{T}+\ldots)$ $(\mathbf{H}_{M}(1)+\Delta_{\mathbf{H}_{M}(1)})=(z+z^{(1)}+\ldots)(\boldsymbol{p}^{T}+(\boldsymbol{p}^{(1)})^{T}+\ldots)(\mathbf{H}_{M}(0)+\Delta_{{\mathbf{H}_{M}(0)}}),$
$(\mathbf{H}_{M}(1)+\Delta_{\mathbf{H}_{M}(1)}))$ $(\boldsymbol{q}+\boldsymbol{q}^{(1)}+\ldots)=(z+z^{(1)}+\ldots)(\mathbf{H}_{M}(0)+\Delta_{\mathbf{H}_{M}(0)}))(\boldsymbol{q}+\boldsymbol{q}^{(1)}+\ldots),$
@@
$|z^{(1)}|$ $=\left|\frac{\boldsymbol{p}^{T}\,(\Delta_{\mathbf{H}_{M}(1)}-z\,\Delta_{\mathbf{H}_{M}(0)})\,\boldsymbol{q}_{L}}{\boldsymbol{p}^{T}\,\mathbf{H}_{M}(0)\,\boldsymbol{q}}\right|$
@@
$\boldsymbol{q}^{(j)}=\boldsymbol{p}^{(j)}=\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{\dagger}\,\boldsymbol{e}_{j},$ (3.5)
@@
$(\boldsymbol{p}^{(j)})^{T}\,\mathbf{H}_{M}(0)\,\boldsymbol{q}^{(j)}$ $=\left(\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{\dagger}\,\boldsymbol{e}_{j}\right)^{T}\,\mathbf{H}_{M}(0)\,\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{\dagger}\,\boldsymbol{e}_{j}$
@@
The ill-disposedness of the eigenvalue $z_{j}$ with respect to unstructured perturbations can be measured by checking whether $\rho_{j}$ is "large". Applying formulas (2.4), (2.5) and (3.5) to (3.7), we obtain for $j=1,\ldots,M$
$\rho_{j}$ $=\frac{1}{|\gamma_{j}|}\left\|\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{\dagger}\,\boldsymbol{e}_{j}\right\|_{2}^{2}\left(|z_{j}|\left\|\mathbf{V}_{M}(\boldsymbol{z})\,\mathbf{\Gamma}_{M}\,\mathbf{V}^{T}_{M}(\boldsymbol{z})\right\|_{2}+\left\|\mathbf{V}_{M}(\boldsymbol{z})\,\mathbf{Z}_{M}\,\mathbf{\Gamma}_{M}\,\mathbf{V}^{T}_{M}(\boldsymbol{z})\right\|_{2}\right).$ (3.8)
@@
where $\kappa(\mathbf{X}):=\|\mathbf{X}\|_{2}\,\|\mathbf{X}^{\dagger}\|_{2}$ denotes the condition number of the matrix $\mathbf{X}$. For diagonal matrices $\mathbf{\Gamma}_{M}$ and $\mathbf{Z}_{M}\,\mathbf{\Gamma}_{M}$, it holds
@@
$\rho_{j}\leq\zeta_{j}\bigg{\|}\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{\dagger}\,\boldsymbol{e}_{j}\bigg{\|}_{2}^{2}\,\big{\|}\mathbf{V}_{M}(\boldsymbol{z})\big{\|}_{2}^{2}.$
@@
$\sum\limits_{j=1}^{M}\rho_{j}\leq\zeta_{\max}\big{\|}\mathbf{V}_{M}(\boldsymbol{z})\big{\|}_{2}^{2}\,\sum\limits_{j=1}^{M}\bigg{\|}\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{\dagger}\,\boldsymbol{e}_{j}\bigg{\|}_{2}^{2}$
@@
$\|\boldsymbol{\rho}\|_{1}\leq\zeta_{\max}\big{\|}\mathbf{V}_{M}(\boldsymbol{z})\big{\|}_{2}^{2}\,\big{\|}\mathbf{V}^{\dagger}_{M}(\boldsymbol{z})\big{\|}_{\mathrm{F}}^{2}$
@@ 3.2 Sensitivity Hankel pencil eigenvalues with structured perturbations
$S^{(j)}_{m}=$ $\frac{1}{\gamma_{j}}\bigg{(}\,\sum\limits_{k=\max\{1,m-M\}}^{\min\{M,m-1\}}\bigg{(}\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{\dagger}\bigg{)}_{k,j}\,\bigg{(}\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{\dagger}\bigg{)}_{m-k,j}$
$-z_{j}\,\sum\limits_{k=\max\{1,m-M+1\}}^{\min\{M,m\}}\bigg{(}\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{\dagger}\bigg{)}_{k,j}\,\bigg{(}\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{\dagger}\bigg{)}_{m-k+1,j}\bigg{)}.$

### v2

@@ 1 Introduction
We study the question of recovery of rational functions (1.1) using as input data a finite set of their Fourier coefficients $\widehat{r}_{k}$ defined by (4.1). We use the property that Fourier coefficients of rational functions have the exponential structure and we reconstruct this exponential structure in the frequency domain applying the ESPRIT method [19, 17]. The idea of recovery of rational functions by estimation of signal parameters in the frequency domain is not new. Similar technique for the recovery of or approximation by rational functions was employed in [7, 20, 21, 22]. The paper [21] is of special interest to us. In [21], the author considered recovery of the rational function (1.1) based on using noisy input data and discussed reconstruction results for the poles $z_{j}$, $j=1,...,M$. It was pointed out that for different levels of noise, not all poles $z_{j}$ are reconstructed with the same accuracy, and sometimes even the locations of certain poles are not reconstructed correctly. From these numerical observations, we can conclude that some poles $z_{j}$ of the rational function (1.1) are more ”sensitive“ to perturbations in the input data than the others. Our goal in this paper is to present theoretical explanations of these numerical observations by applying the sensitivity analysis. The method developed in [21] employees the AAK theory [16] for the reconstruction of the exponential structure of the Fourier coefficients $\widehat{r}_{k}$. We modify these ideas, by applying the ESPRIT method in the frequency domain. Since ESPRIT is based on using the Hankel pencil, it makes our method suitable for the sensitivity analysis.
@@
In [23], the authors developed the sensitivity analysis for the eigenvalues of Loewner pencil with respect to the unstructured and structured perturbations. We referee the interested reader to [1] for further application of Loewner framework for model reduction of dynamical systems. Following approach in [23], we present the sensitivity analysis for the poles of $r$ in (1.1) computed by our method. According to our idea, poles located inside and outside the unit circle are reconstructed separately as eigenvalues of some Hankel matrix pencils. Further, we consider two kinds of Hankel pencil perturbations. For the sensitivity $\rho$ with respect to unstructured perturbations, we present the connection with the condition numbers of Vandermonde matrices that are used in factorizations of the corresponding Hankel matrices. The sensitivity $\eta$ with respect to structured perturbations of Hankel pencil is studied in [24]. We present modification of these results for our special rational reconstruction problem (1.1). Note that the theory of sensitivities of eigenvalues of Hankel pencil perturbations was considered also in [11, 3]. In [11] and [3], the authors measured both structured and unstructured perturbations by the same quantity. Similarly as in [23], we introduce two different quantities for structured $\eta$ and unstructured $\rho$ sensitivities, respectively, of poles of the rational function (1.1). By checking if $\rho$ or $\eta$ is ”large“ for some pole $z_{j}$, we measure the ill-disposedness of $z_{j}$, while consideration of $\eta$ gives us also the additional information regarding that to perturbation of which measurement $\widehat{r}_{k}$ the pole $z_{j}$ is the most sensitive.
@@
The method developed in this paper is also closely related to our ESPIRA method for the exponential recovery [8][10]. ESPIRA uses the property that the classical Fourier coefficients of exponential sums have a rational structure and reconstructs this rational structure with the AAA method for rational approximation [14] in the frequency domain. Thus, the idea of rational reconstruction used in this paper can be considered as converse to the idea of ESPIRA. We want to pay a special attention also to the paper [5], where authors developed a method for exponential reconstruction using the sensitivity analysis of Hankel pencil eigenvalues.
Finally, we want to mention the other methods for the recovery of rational functions such as the AAA method [14], the RKFIT (Rational Krylov Fitting) method [4], the Vector Fitting algorithm [13] and the Loewner framework method [1]. In [12], the authors discussed all these algorithms for the recovery of rational functions using as input information given noisy set of samples of $r$. Since we are interested in the sensitivity analysis of the poles, we have to assume that we apply perturbations to the Fourier coefficients $\widehat{r}_{k}$. Note that we can introduce the corresponding modification of the method based on employing function value of $r$ by computing $\widehat{r}_{k}$ via the quadrature formula for some grid of points (see Remark 4.1). Modifying the algorithm in this way will allow us to compare its performance with known rational methods. Numerical experiments show us that our approach gives equally good and for some examples even better recovery results from given noisy function values than the AAA method.
@@
Outline. In Section 2, we present the main ideas of the ESPRIT method for the exponential recovery. In Section 3, we consider the sensitivity analysis for Hankel pencil eigenvalues: Subsection 3.1 is dedicated to unstructured perturbations and in Subsection 3.2, we consider structured perturbations. In Section 4, we describe our main results. In Subsection 4.1, we present main ideas of our new approach for the recovery of rational functions (1.1). Subsection 4.2 is dedicated to the sensitivity analysis of poles reconstructed with our method. Finally, in Section 5, we present numerical experiments.
@@
Notation. Let us agree that throughout the paper, we will use the matrix notation $\mathbf{X}_{L,K}:=(X_{ij})_{i,j=1}^{L,K}$ for complex matrices of size $L\times K$ and the submatrix notation $\mathbf{X}_{L,K}(m:n,k:\ell)$, to denote a submatrix of $\mathbf{X}_{L,K}$ with rows indexed $m$ to $n$ and columns indexed $k$ to $\ell$, where first row and first column has index 1 (even though the row and column indices for the definition of the matrix may start with 0). For square matrices, we often use the short notation $\mathbf{X}_{K}$ instead of $\mathbf{X}_{K,K}$. By $\mathbf{X}_{L,K}^{\dagger}$ we denote the Moore-Penrose inverse of $\mathbf{X}_{L,K}$ and by $\mathbf{X}_{K}^{-1}$ inverse of the invertible matrix $\mathbf{X}_{K}$. We consider $\|\mathbf{X}_{L,K}\|_{2}:=\sigma_{\max}(\mathbf{X}_{L,K})$ the 2-norm or spectral norm, and $\|\mathbf{X}_{L,K}\|_{\mathrm{F}}:=\left(\sum\limits_{i=1}^{L}\sum\limits_{j=1}^{K}|X_{ij}|^{2}\right)^{1/2}=\left(\sum\limits_{i=1}^{\min\{L,K\}}\sigma_{i}^{2}\right)^{1/2}$ the Frobenius norm of the matrix $\mathbf{X}_{L,K}$. The $\ell_{1}$- and $\ell_{2}$-norm of the vector $\boldsymbol{x}=(x_{1},\ldots,x_{K})$ are given as usually by $\|\boldsymbol{x}\|_{1}=\sum\limits_{j=1}^{K}|x_{j}|$ and $\|\boldsymbol{x}\|_{2}=\left(\sum\limits_{j=1}^{K}|x_{j}|^{2}\right)^{1/2}$ respectively.
@@ 3 Sensitivities of Hankel pencil eigenvalues
The concept of sensitivities arises from the perturbation of eigenvalues of the matrix pencil and for Hankel matrices was introduced in [11] and further considered in [3]. In [23], the notion of sensitivities was defined for Loewner matrix pencil. In the paper [23], the authors define two sensitivities, which correspond to structured and unstructured perturbations. Using similar ideas, in [24], the author considered the structured sensitivity of Hankel matrix pencil. In [5], the sensitivity analysis of Hankel pencil was applied to develop a new method for exponential analysis for the recovery from noisy data. In this section, we consider sensitivities of eigenvalues of the Hankel pencil (2.3) and further in Sunbsection 4.2 we present modification of the results from this section to develop sensitivity analysis for the poles of rational functions.
@@ 3.1 Sensitivity of Hankel pencil eigenvalues with unstructured perturbations
$(\boldsymbol{p}^{T}+(\tilde{\boldsymbol{p}}^{(1)})^{T}+\ldots)$ $(\mathbf{H}_{M}(1)+\Delta_{\mathbf{H}_{M}(1)})=(z+z^{(1)}+\ldots)(\boldsymbol{p}^{T}+(\tilde{\boldsymbol{p}}^{(1)})^{T}+\ldots)(\mathbf{H}_{M}(0)+\Delta_{{\mathbf{H}_{M}(0)}}),$
$(\mathbf{H}_{M}(1)+\Delta_{\mathbf{H}_{M}(1)}))$ $(\boldsymbol{q}+\tilde{\boldsymbol{q}}^{(1)}+\ldots)=(z+z^{(1)}+\ldots)(\mathbf{H}_{M}(0)+\Delta_{\mathbf{H}_{M}(0)}))(\boldsymbol{q}+\tilde{\boldsymbol{q}}^{(1)}+\ldots),$
@@
$|z^{(1)}|$ $=\left|\frac{\boldsymbol{p}^{T}\,(\Delta_{\mathbf{H}_{M}(1)}-z\,\Delta_{\mathbf{H}_{M}(0)})\,\boldsymbol{q}}{\boldsymbol{p}^{T}\,\mathbf{H}_{M}(0)\,\boldsymbol{q}}\right|$
@@
$\boldsymbol{q}^{(j)}=\boldsymbol{p}^{(j)}=\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{-1}\,\boldsymbol{e}_{j},$ (3.5)
@@
$(\boldsymbol{p}^{(j)})^{T}\,\mathbf{H}_{M}(0)\,\boldsymbol{q}^{(j)}$ $=\left(\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{-1}\,\boldsymbol{e}_{j}\right)^{T}\,\mathbf{H}_{M}(0)\,\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{-1}\,\boldsymbol{e}_{j}$
@@
The ill-disposedness of the eigenvalue $z_{j}$ with respect to unstructured perturbations can be measured by checking whether $\rho_{j}$ is ”large“. Applying formulas (2.4), (2.5) and (3.5) to (3.7), we obtain for $j=1,\ldots,M$
$\rho_{j}$ $=\frac{1}{|\gamma_{j}|}\left\|\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{-1}\,\boldsymbol{e}_{j}\right\|_{2}^{2}\left(|z_{j}|\left\|\mathbf{V}_{M}(\boldsymbol{z})\,\mathbf{\Gamma}_{M}\,\mathbf{V}^{T}_{M}(\boldsymbol{z})\right\|_{2}+\left\|\mathbf{V}_{M}(\boldsymbol{z})\,\mathbf{Z}_{M}\,\mathbf{\Gamma}_{M}\,\mathbf{V}^{T}_{M}(\boldsymbol{z})\right\|_{2}\right).$ (3.8)
@@
where $\kappa(\mathbf{X}):=\|\mathbf{X}\|_{2}\,\|\mathbf{X}^{-1}\|_{2}$ denotes the condition number of the matrix $\mathbf{X}$. For diagonal matrices $\mathbf{\Gamma}_{M}$ and $\mathbf{Z}_{M}\,\mathbf{\Gamma}_{M}$, it holds
@@
Note that we can proceed further with the estimation of (3.10) applying known results for upper bounds of condition numbers of the Vandermonde matrices [15, 2].
@@
$\rho_{j}\leq\zeta_{j}\bigg{\|}\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{-1}\,\boldsymbol{e}_{j}\bigg{\|}_{2}^{2}\,\big{\|}\mathbf{V}_{M}(\boldsymbol{z})\big{\|}_{2}^{2}.$
@@
$\sum\limits_{j=1}^{M}\rho_{j}\leq\zeta_{\max}\big{\|}\mathbf{V}_{M}(\boldsymbol{z})\big{\|}_{2}^{2}\,\sum\limits_{j=1}^{M}\bigg{\|}\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{-1}\,\boldsymbol{e}_{j}\bigg{\|}_{2}^{2}$
@@
$\|\boldsymbol{\rho}\|_{1}\leq\zeta_{\max}\big{\|}\mathbf{V}_{M}(\boldsymbol{z})\big{\|}_{2}^{2}\,\big{\|}\mathbf{V}^{-1}_{M}(\boldsymbol{z})\big{\|}_{\mathrm{F}}^{2}$
@@ 3.2 Sensitivity Hankel pencil eigenvalues with structured perturbations
$S^{(j)}_{m}=$ $\frac{1}{\gamma_{j}}\bigg{(}\,\sum\limits_{k=\max\{1,m-M\}}^{\min\{M,m-1\}}\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{-1}\bigg{)}_{k,j}\,\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{-1}\bigg{)}_{m-k,j}$
$-z_{j}\,\sum\limits_{k=\max\{1,m-M+1\}}^{\min\{M,m\}}\bigg{(}\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{-1}\bigg{)}_{k,j}\,\bigg{(}\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{-1}\bigg{)}_{m-k+1,j}\bigg{)}.$